OpenStudy (anonymous):

Find the quotient of the following complex numbers x=5(cos15 + isin15) y= 3(cos70 + isin70) Express your answer in trigonometric form.

OpenStudy (jdoe0001):

\(\large \bf \cfrac{r_1[cos(\alpha)+isin(\alpha)]}{r_2[cos(\beta)+isin(\beta)]}\implies \cfrac{r_1}{r_2}[cos(\alpha - \beta)+isin(\alpha - \beta)] \)

OpenStudy (anonymous):

Okay... so it would be 5/3(cos305 + isin305)?

OpenStudy (jdoe0001):

5/3 is right.. .... 15 - 70 = 305?

OpenStudy (jdoe0001):

hmmm ohh.. I see.. you're using the positive angle... then yes

OpenStudy (anonymous):

My thanks

OpenStudy (jdoe0001):

yw